A force parallel to the π‘₯-axis acts on a particle moving along the π‘₯-axis. This force produces potential energy π‘ˆ(π‘₯) given by π‘ˆ(π‘₯) = 𝛼π‘₯ 4 , where 𝛼 = 0.630 J/m4 . What is the force (magnitude and direction) when the particle is at π‘₯ = βˆ’0.800 m?


1
Expert's answer
2021-10-27T08:51:40-0400

Given:

"U(x)=\alpha x^4"

"\alpha=0.630 \:\rm J/m^4"

"x=-0.800\:\rm m"


The force

"F=-\frac{dU}{dx}=-4\alpha x^3"

"F=-4*0.800*(-0.800)^3=1.64\:\rm N"